OpenStudy (anonymous):

Quick help please? Which system of equations matches the graph below?...Is my answer correct?

OpenStudy (anonymous):

Answers are: A. y = x2 – 1 if x > –1 y = 2x – 4 if x –2 B. y = x2 – 1 if x < 1 y = 2x – 4 if x 1 C. y = x2 – 1 if x < 0 y = 2x – 4 if x –2 D. y = x2 – 1 if x < 1 y = 2x – 4 if x –2 I got...B.

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

thank you! ur the first person that has answered and Ive had it posted for 40 mins.